OpenStudy (anonymous):

Write the polynomials in factored form.? 1.4x^3+8x^2-96x 2.x^3+9x^2+18x Please explain and show work.

OpenStudy (anonymous):

in order to write the polynomial in standard form have to suuply coefficients of missing powers of x but i think urs is already in standard form if not just add +0 to the right of the question

OpenStudy (anonymous):

I meant factored

OpenStudy (anonymous):

oops 1) 4x^3+8x^2-96x=4x(x^2 +2x -24) =4x(x^2+6x -4x -24)=4x(x(x+6)-4(x+6))=4x(x-4)(x+6)

OpenStudy (anonymous):

x^3+9x^2+18x=x(x^2 +9x +18)=x(x^2+6x+3x+18) =x(x(x+6) +3(x+6))=x(x+3)(x+6)

OpenStudy (anonymous):

(4x^3+8x^2-96x) : 4\[x^3+2x^2-24\]

OpenStudy (anonymous):

then factor

OpenStudy (anonymous):

soory

OpenStudy (anonymous):

x^3+2x^2-24x

OpenStudy (anonymous):

x(x^2+2x-24) x(x-4)(x+6)